Expand in Taylor series at both sides to find the error . Sa+y (3) ds = (34 (2) – y (x – h)) + error

dont use matlab
Answer #1

ANSWER:-

Expand in Taylor series at both side to find the error

By fundamental then On internal calculus

Y(t) = Ft) for some func

\therefore \int_{x}^{x+h}Y(s)\,\,ds

=\int_{x}^{x+h}F'(s)\,\,ds

=F(x+h)-F(x)

=F(x)+h\, F'(x)+\frac{h^2}{2!}F"(x) for some C_1 between T and x+h

= h\, y(x) +\frac{h^{2}}{2!}y'(x) +\frac{h^{3}}{3!}y''c_{1}.......... (1)as y(t) =F'(t) \, \forall\, t

Also, by Tylor series expansion

=\frac{h}{2}\left ( 3 y(x) -y(x-h) \right )

=\frac{h}{2} \left [3 y(x) -\left \{ y(x) -hy'(x) +\frac{h^{2}}{2!}y"(c_{2})\right \} \right ]

For some C_2 between x-h and x

=\frac{h}{2} \left [2y(x) +hy'(x) -\frac{h^{2}}{2!}y"(c_{2}) \right ]

=h \, y(x)+\frac{h^{2}}{2!}y'(x) -\frac{h^{3}}{2.2!}y"(c_{2}) .......... (2)

Therefore from equation (1) and (2)

=\int_{x}^{x+h}y(s) ds-\frac{h}{2}\left [ 3y(x) -y(x-h) \right ]

=h^{3}\left ( \frac{y"(c_{1})}{3!}-\frac{y"(c_{2})}{2.2!} \right )

=0(h^{3})

Therefore The error is of 0(h^{3})